# Wave propagation in an elastic lattice with non-reciprocal stiffness and engineered damping

### v1

@@ III Lattice with non-reciprocal damping
This section considers a system with reciprocal stiffness ($\alpha=0$) and non-reciprocal intersite damping. The damping in the forward direction is taken to be $c_{f}=c(1+\beta)$, while in the backward direction it is $c_{b}=c(1-\beta)$, where $\beta\in[-1,1]$. The governing equation of motion for this model, ignoring onsite damping, is
$m\ddot{u}_{n}+c(2\dot{u}_{n}-(1+\beta)\dot{u}_{n-1}-(1-\beta)\dot{u}_{n+1})\\ +(2k+k_{g})u_{n}-k(u_{n-1}+u_{n+1})=0.$ (23)
@@
These results reveal a different non-reciprocal mechanism compared to the case of non-reciprocal springs. The presence of the term proportional to $\sin q$ in Eq. (25) makes the oscillation frequency $\text{Re}[\omega(q)]$ also asymmetric with respect to $q=0$.
For small damping ($c\sim\mathcal{O}(\epsilon)$) and non-reciprocity ($\beta\sim\mathcal{O}(\epsilon)$), the approximate dispersion relations are:
$\text{Re}[\omega(q)]$ $=\sqrt{A_{0}(q)}-\frac{c^{2}(1-\cos q)^{2}}{2m^{2}\sqrt{A_{0}(q)}}+\frac{c\beta}{m}\sin q+\mathcal{O}(\epsilon^{3}),$ (27)
$\text{Im}[\omega(q)]$ $=-\frac{c}{m}(1-\cos q)+\mathcal{O}(\epsilon^{3}),$ (28)
where $A_{0}(q)=\frac{1}{m}[k_{g}+2k(1-\cos q)]$. The leading-order terms indicate that the real component has an asymmetric shift proportional to $\beta$, while the imaginary component is nearly symmetric but negative, implying that the wave will attenuate in both directions at approximately the same rate.
@@
Special non-reciprocal damping: Gyroscopic damping
The preceding model exhibits both an asymmetric real dispersion relation and global wave decay. To isolate the frequency asymmetry from the dissipative effects, we now define a special non-reciprocal damper that leads to purely oscillatory dynamics. As depicted in Fig. 7(a), this corresponds to an effective damping of $c\beta$ in the forward direction and $-c\beta$ in the backward direction. This is akin to gyroscopic damping that is not dissipative Carta et al. (2014); Baz (2020); Attarzadeh et al. (2019); Wang et al. (2015). This construction isolates the non-reciprocal behavior from the net dissipation inherent in standard intersite damping. In addition, this special case will later be combined with non-reciprocal stiffness to explore novel wave dynamics.
@@
The group velocity consequently loses its symmetry as well. Figure 7(c) shows a parametric plot demonstrating that, in contrast to the case of non-reciprocal stiffness, points of zero group velocity can emerge at wavenumbers other than $q=0$ or $q=\pm\pi$. Interestingly, a similar observation is made related to the modification of Brillouin zone boundaries in the context of non-reciprocal Willis monatomic lattices Al Ba’ba’a (2023).
@@
$\beta\approx{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-}\frac{\sqrt{m}k}{c\sqrt{k_{g}}}q.$ (32)
@@
Figure 8: Numerical simulation of wave dynamics in a 1D lattice with gyroscopic damping. (a) Space-time colormap of the displacement field $u_{n}(t)$ in a 200-particle finite lattice with fixed boundaries, illustrating symmetric wave amplitudes but asymmetric propagation speeds for left- and right-travelling wave packets. (b) Time evolution of the instantaneous global maximum, showing no net amplification or decay. (c–d) Velocity responses and their spectral amplitudes at sites $n=40$ and 160160160 reveal the coexistence of two wave components with distinct group velocities, corresponding to slightly different peak frequencies ($\omega\approx 23.00$ and $\omega\approx 25.75$).
@@
As predicted by the purely real dispersion relation [Eq. (30)], the system is stable and exhibits no exponential growth or decay. This stability is corroborated by the global displacement norm, which remains bounded throughout the simulation [Fig. 8(b)]. The velocity time series at two symmetric probe locations ($n=40$ and $n=160$), shown in Fig. 8(c), confirms the different arrival times and shows that the wave packets maintain comparable amplitudes. A key finding is the emergence of spectral asymmetry: the velocity spectra in Fig. 8(d) reveal distinct peak frequencies for the forward-propagating ($\omega\approx 25.75$) and backward-propagating ($\omega\approx 23.00$) waves. This direction-dependent frequency shift is a direct consequence of the asymmetric dispersion shown in Fig. 7(b).
These finite-lattice simulations, therefore, confirm that this gyroscopic damper introduces direction-dependent group velocities and frequency shifts without inducing exponential amplification or attenuation.
@@
IV Lattice with Combined Non-reciprocal Stiffness and Damping
@@
$m\ddot{u}_{n}+c_{g}\dot{u}_{n}+c\beta(\dot{u}_{n+1}-\dot{u}_{n-1})+(2k+k_{g})u_{n}\\ -k(1+\alpha)u_{n-1}-k(1-\alpha)u_{n+1}=0.$ (33)
@@
These results reveal a remarkable decoupling of the two non-reciprocal effects at leading order. The imaginary part of the frequency, which governs temporal growth or decay, is determined entirely by the non-reciprocal stiffness parameter $\alpha$, and matches the expression for lattice with nonreciprocal stiffness and onsite damping. The real part, on the other hand, acquires an asymmetric contribution from the non-reciprocal damper $\beta$, inducing a direction-dependent frequency shift without altering the system’s gain or loss characteristics. This enables independent tuning of wave amplification and oscillation frequency through parameters $\alpha$ and $\beta$, respectively.
Figures 9(b) and 9(c) present the real and imaginary parts of the complex dispersion relation, $\omega(q)$, for a fixed stiffness asymmetry ($\alpha=0.1$) to investigate the influence of the parameter $\beta$. While the imaginary part, $\text{Im}[\omega(q)]$, is nearly identical to the case where $\beta=0$, the real part, $\text{Re}[\omega(q)]$, becomes asymmetric for non-zero values of $\beta$.
This behavior is further detailed in the parametric maps shown in Figs. 9(d) and 9(e). The group velocity map [Fig. 9(d)] exhibits a strong asymmetry in both sign and magnitude that is tunable by $\beta$. In contrast, the growth rate map [Fig. 9(e)] confirms that $\text{Im}[\omega(q)]$ is largely insensitive to $\beta$. The inherent asymmetry in the growth rate with respect to the wavenumber $q$ is therefore preserved, as it is governed by the nonreciprocal stiffness $\alpha$.
Figure 10: Finite lattice simulation illustrating amplification dependence on $\beta$ in a system with combined nonreciprocal stiffness and gyroscopic damping. (a,b) Space–time plots of normalized displacement for $\beta=\pm 0.2$. (c,d) Corresponding time-domain velocity responses at $n=130,160$. The case with $\beta=0.2$ shows a slightly more amplified wavepacket due to slower group velocity. (e,f) Velocity spectra reveal dominant frequencies. We take $\alpha=0.1$ and keep other parameters the same as before.
@@
The results reveal that the sign of $\beta$ provides a mechanism to control wave dynamics. Figures 10(a) and 10(b) show that the forward-propagating wave is amplified in both cases, a characteristic governed by the nonreciprocal stiffness $\alpha$. However, switching the sign of $\beta$ alters the group velocity of the wave packet. This is explicitly confirmed by the velocity time series at downstream sites ($n=130$ and $n=160$) in Figs. 10(c) and (d), which show that the wave packet for $\beta=-0.2$ arrives at a given site faster than for $\beta=+0.2$.
@@
Interestingly, while the amplification rate is set by $\alpha$, the net amplification at a specific site depends on the group velocity as well. For $\beta=+0.2$, the wave packet travels slower, allowing more time for the amplification, resulting in a larger amplitude at a given location compared to the $\beta=-0.2$ case. The net amplification is therefore a combined effect of the growth rate (dictated by $\alpha$) and the propagation time (dictated by $\beta$). This is one of the key results of our work.
Finally, Figs. 10(e) and 10(f) show that the sign of $\beta$ also shifts the dominant frequency of the amplified wave, along with a gain in the spectral content. This tunability is further explored in Fig. 11, which plots the dominant amplified frequency as a function of $\alpha$ and $\beta$ for a forward-propagating wave. The frequency exhibits a nearly quadratic dependence on $\alpha$ for a fixed $\beta$ [Fig. 11(a)], consistent with the higher-order corrections in Eq. (12). In contrast, Fig. 11(b) demonstrates that varying $\beta$ provides an effective route to tune the wave’s frequency while leaving the amplification rate, governed by $\alpha$, nearly unchanged.
In summary, while the amplification rate of waves is primarily determined by $\alpha$, the sign and magnitude of $\beta$ can be used to tune net amplification (through changing the group velocity) and the dominant frequency of the wave packet.
@@
Figure 11: Effect of $\alpha$ and $\beta$ on dominant oscillation frequency and temporal growth at $q=0.6\pi$. (a) Real part of the dominant eigenfrequency $\text{Re}(\omega)$ plotted as a function of stiffness asymmetry $\alpha$ for $\beta=-0.2,0,0.2$. Change in $\beta$ leads to a significant change in oscillation frequency for a fixed $\alpha$. (b) Imaginary part $\text{Im}(\omega)$ for the same parameter sweep shows minimal variation, suggesting that damping asymmetry has a negligible effect on the temporal growth rate. Solid lines represent analytical results from the complex dispersion relation, and circular markers indicate numerically extracted peak frequencies from finite-lattice simulations (see Fig. 10). The inset in (b) confirms close agreement between analytical and numerical growth rates at $\alpha=0.1$.
Figure 12: Effect of boundaries on wave mixing and amplification for $|q^{*}|=0.6\pi$. (a,b) Ray diagram for the evolution of the initial wave packet placed near the left boundary. Backward traveling wave has a different group velocity, and then it reflects to mix with the forward amplifying wave. (c,d) Spatiotemporal velocity waterfall plots, illustrating wave propagation and boundary interactions as one changes $\beta$. The backward-propagating wave reflects off the left boundary and re-enters the lattice, leading to additional interaction with the medium. (e,f) Velocity response at site 80 captures both the initial forward-propagating wave and the subsequent reflected wave packet. (g,h) Frequency spectra at site 80 display two distinct peaks associated with the forward and reflected modes but in different relative amplitudes.
Wave mixing through boundary reflection
@@
In a finite system, boundaries are not merely passive reflectors but can be actively utilized to manipulate wave properties. Here, we investigate how boundary reflection can induce “wave mixing”, providing an additional mechanism for tuning the amplified signal’s frequency content. We achieve this by launching a wave packet near one end of the lattice. The backward-propagating (attenuated) component travels to the opposite boundary, reflects, and then propagates forward, mixing with the original forward-propagating (amplified) component. As we will show, the nonreciprocal damping parameter $\beta$ is crucial in controlling the effectiveness of this mixing. Figure 12 compares the dynamics for $\beta=-0.2$ (left column) and $\beta=+0.2$ (right column), keeping the stiffness asymmetry fixed at $\alpha=0.1$.
The ray diagram in Figs. 12(a) and 12(b) illustrate this process. The backward-propagating wave has a significantly faster group velocity for $\beta=+0.2$ compared to $\beta=-0.2$. Consequently, it reaches the boundary and reflects more quickly, experiencing less attenuation during its transit. This difference in travel time and attenuation directly impacts the local dynamics, as seen in the velocity waterfall plots in Figs. 12(c) and 12(d). When we extract velocity time series at site $n=80$ [Figs. 12(e) and 12(f)] for these cases, the difference is clear. For $\beta=-0.2$, the slower backward wave is heavily attenuated, resulting in a reflected packet that is well-separated in time and has a modest amplitude. For $\beta=+0.2$, the faster, less-attenuated backward wave produces a much stronger reflected packet that arrives sooner, enabling a more effective mixing of the two components.
@@
The consequence of this wave mixing is evident in the frequency spectra of the signal at site $n=80$, shown in Figs. 12(g) and 12(h). Recall that the forward and backward waves have different dominant frequencies. For $\beta=+0.2$, the effective mixing of the strong reflected wave with the forward wave results in a spectrum where two distinct frequency peaks, $\omega\approx 25.86$ (forward wave) and $\omega\approx 23.00$ (reflected wave) are prominent and significantly amplified. In contrast, for $\beta=-0.2$, the spectrum is dominated by the forward wave ($\omega\approx 22.89$) and reflected wave ($\omega\approx 25.63$), but the weak reflection has a smaller impact. The value of $\beta$ along with the boundary thus acts as a switch, controlling which frequencies are mixed and their relative dominance. Therefore, these results demonstrate that boundaries in nonreciprocal lattices can be seen as functional elements that enable complex wave-shaping phenomena in addition to wave amplification.

### v2

@@ III Lattice with non-reciprocal damping
This section considers a system with reciprocal stiffness ($\alpha=0$) and non-reciprocal intersite damping as shown in Fig. 8a. The damping in the forward direction is taken to be $c_{f}=c(1+\beta)$, while in the backward direction it is $c_{b}=c(1-\beta)$, where $\beta\in[-1,1]$. The governing equation of motion for this model, ignoring onsite damping, is
$\begin{split}m\ddot{u}_{n}&+c(2\dot{u}_{n}-(1+\beta)\dot{u}_{n-1}-(1-\beta)\dot{u}_{n+1})\\ &+(2k+k_{g})u_{n}-k(u_{n-1}+u_{n+1})=0.\end{split}$ (26)
@@
These results reveal a different non-reciprocal mechanism compared to the case of non-reciprocal springs. The presence of the term proportional to $\sin q$ in Eq. (28) makes the oscillation frequency $\text{Re}[\omega(q)]$ also asymmetric with respect to $q=0$. This spectral asymmetry implies that forward- and backward-propagating waves (at wavenumbers $\pm q$) are characterized by distinct oscillation frequencies and group velocities. Moreover, the imaginary component $\text{Im}[\omega(q)]$ also acquires an asymmetric profile for nonzero $\beta$, as illustrated in the dispersion curves in Figs. 8(b) and 8(c). This indicates that the nonreciprocity parameter $\beta$ simultaneously dictates both the wave kinematics (speed/frequency) and the temporal envelope evolution (growth/decay rates).
While this general model demonstrates the broad impact of nonreciprocal damping, it highlights a limitation regarding control: $\beta$ introduces asymmetry in both the real and imaginary dispersion curves concurrently. This coupling prevents the independent manipulation of signal timing (velocity) and signal magnitude (dissipation/gain). To isolate the kinematic nonreciprocity from the dissipative effects, we next examine a specialized form of nonreciprocal damping inspired by gyroscopic mechanics.
Figure 8: Wave propagation in a 1D lattice with nonreciprocal damping. (a) Schematic of the spring–mass chain with direction-dependent damping coefficients $c_{f}=c(1+\beta)$ and $c_{b}=c(1-\beta)$, along with reciprocal intersite stiffness $k$ and onsite stiffness $k_{g}$. (b) Real and (c) imaginary parts of the dispersion relation $\omega(q)$ for reciprocal ($\beta=0$) and nonreciprocal ($\beta=0.2$) cases. The real and imaginary parts exhibit asymmetry. The imaginary part is predominantly negative, indicating dissipation. We take $c=7.5$ and keep other parameters identical to the reciprocal case.
Special non-reciprocal damping: Gyroscopic damping
We define a special non-reciprocal damper that leads to purely oscillatory dynamics. As depicted in Fig. 9(a), this corresponds to an effective damping of $c\beta$ in the forward direction and $c\beta$ in the backward direction. This is akin to gyroscopic damping that is not dissipative Carta et al. (2014); Baz (2020); Attarzadeh et al. (2019); Wang et al. (2015). This construction isolates the non-reciprocal behavior from the net dissipation inherent in standard intersite damping. In addition, this special case will later be combined with non-reciprocal stiffness to explore novel wave dynamics.
@@
The group velocity consequently loses its symmetry as well. Figure 9(c) shows a parametric plot demonstrating that, in contrast to the case of non-reciprocal stiffness, points of zero group velocity can emerge at wavenumbers other than $q=0$ or $q=\pm\pi$. Interestingly, a similar observation is made related to the modification of Brillouin zone boundaries in a physically distinct system of a translating elastic rod Al Ba’ba’a (2023). In that work, a momentum bias introduces a velocity-dependent coupling term in the discrete framework, which is mathematically analogous to the gyroscopic damping presented in our study.
@@
$\beta\approx{-}\frac{\sqrt{m}k}{c\sqrt{k_{g}}}q.$ (33)
@@
Figure 10: Finite-chain simulations of wave dynamics in a 1D lattice with gyroscopic damping. (a) Space-time colormap of the displacement field $u_{n}(t)$ in a 200-particle finite lattice with fixed boundaries, illustrating symmetric wave amplitudes but asymmetric propagation speeds for left- and right-travelling wave packets. (b) Time evolution of the instantaneous global maximum, showing no net amplification or decay. (c–d) Velocity responses and their spectral amplitudes at sites $n=40$ and $160$ reveal the coexistence of two wave components with distinct group velocities, corresponding to slightly different peak frequencies ($\omega\approx 23.00$ and $\omega\approx 25.75$).
@@
As predicted by the purely real dispersion relation [Eq. (31)], the system is stable and exhibits no exponential growth or decay. This stability is corroborated by the global displacement norm, which remains bounded throughout the simulation [Fig. 10(b)]. The velocity time series at two symmetric probe locations ($n=40$ and $n=160$), shown in Fig. 10(c), confirms the different arrival times and shows that the wave packets maintain comparable amplitudes. A key finding is the emergence of spectral asymmetry: the velocity spectra in Fig. 10(d) reveal distinct peak frequencies for the forward-propagating ($\omega\approx 25.75$) and backward-propagating ($\omega\approx 23.00$) waves. This direction-dependent offset in frequency is a direct consequence of the asymmetric dispersion shown in Fig. 9(b).
These finite-lattice simulations, therefore, confirm that this gyroscopic damper introduces direction-dependent group velocities and offset in frequency without inducing exponential amplification or attenuation.
@@
IV Lattice with Combined Non-reciprocal Stiffness and Damping
@@
$\begin{split}m\ddot{u}_{n}&+c_{g}\dot{u}_{n}+c\beta(\dot{u}_{n+1}-\dot{u}_{n-1})\\ &+(2k+k_{g})u_{n}-k(1+\alpha)u_{n-1}-k(1-\alpha)u_{n+1}=0.\end{split}$ (34)
@@
These results reveal a remarkable decoupling of the two non-reciprocal effects at leading order. The imaginary part of the frequency, which governs temporal growth or decay, is determined entirely by the non-reciprocal stiffness parameter $\alpha$, and matches the expression for lattice with nonreciprocal stiffness and onsite damping. The real part, on the other hand, acquires an asymmetric contribution from the non-reciprocal damper $\beta$, inducing a direction-dependent offset in frequency without altering the system’s gain or loss characteristics. This enables independent tuning of wave amplification and oscillation frequency through parameters $\alpha$ and $\beta$, respectively.
Figures 11(b) and 11(c) present the real and imaginary parts of the complex dispersion relation, $\omega(q)$, for a fixed stiffness asymmetry ($\alpha=0.1$) to investigate the influence of the parameter $\beta$. While the imaginary part, $\text{Im}[\omega(q)]$, is nearly identical to the case where $\beta=0$, the real part, $\text{Re}[\omega(q)]$, becomes asymmetric for non-zero values of $\beta$.
This behavior is further detailed in the parametric maps shown in Figs. 11(d) and 11(e). The group velocity map [Fig. 11(d)] exhibits a strong asymmetry in both sign and magnitude that is tunable by $\beta$. In contrast, the growth rate map [Fig. 11(e)] confirms that $\text{Im}[\omega(q)]$ is largely insensitive to $\beta$. The inherent asymmetry in the growth rate with respect to the wavenumber $q$ is therefore preserved, as it is governed by the nonreciprocal stiffness $\alpha$.
Figure 12: Finite lattice simulation illustrating amplification dependence on $\beta$ in a system with combined nonreciprocal stiffness and gyroscopic damping. (a,b) Space–time plots of normalized displacement for $\beta=\pm 0.2$. (c,d) Corresponding time-domain velocity responses at $n=130,160$. The case with $\beta=0.2$ shows a slightly more amplified wavepacket due to slower group velocity. (e,f) Velocity spectra reveal peak frequencies. We take $\alpha=0.1$ and keep other parameters the same as before.
@@
The results reveal that the sign of $\beta$ provides a mechanism to control wave dynamics. Figures 12(a) and 12(b) show that the forward-propagating wave is amplified in both cases, a characteristic governed by the nonreciprocal stiffness $\alpha$. However, switching the sign of $\beta$ alters the group velocity of the wave packet. This is explicitly confirmed by the velocity time series at downstream sites ($n=130$ and $n=160$) in Figs. 12(c) and (d), which show that the wave packet for $\beta=-0.2$ arrives at a given site faster than for $\beta=+0.2$. This is consistent with the evolution of temporal peak for the forward propagating amplified wave packet (see Fig. 15 in Appendix A).
@@
Interestingly, while the amplification rate is set by $\alpha$, the net amplification at a specific site depends on the group velocity as well. For $\beta=+0.2$, the wave packet travels slower, allowing more time for the amplification, resulting in a larger amplitude at a given location compared to the $\beta=-0.2$ case. The net amplification (seen from the temporal peaks of the Gaussian packets) is therefore a combined effect of the growth rate (dictated by $\alpha$) and the propagation time (dictated by $\beta$). This is one of the key results of our work.
@@
Figure 13: Effect of $\alpha$ and $\beta$ on oscillation frequency and temporal growth at $q=0.6\pi$. (a) $\text{Re}(\omega)$ plotted as a function of stiffness asymmetry $\alpha$ for $\beta=-0.2,0,0.2$. Change in $\beta$ leads to a significant change in oscillation frequency for a fixed $\alpha$. (b) $\text{Im}(\omega)$ for the same parameter sweep shows minimal variation, suggesting that damping asymmetry has a negligible effect on the temporal growth rate. Solid lines represent analytical results from the complex dispersion relation, and circular markers indicate numerically extracted peak frequencies from finite-lattice simulations (see Fig. 12). The inset in (b) confirms close agreement between analytical and numerical growth rates at $\alpha=0.1$.
Finally, Figs. 12(e) and 12(f) show that the sign of $\beta$ also shifts the peak frequency of the amplified wave, along with a gain in the spectral content. This tunability is further explored in Fig. 13, which plots the amplified peak frequency as a function of $\alpha$ and $\beta$ for a forward-propagating wave. The frequency exhibits a nearly quadratic dependence on $\alpha$ for a fixed $\beta$ [Fig. 13(a)], consistent with the higher-order corrections in Eq. (15). In contrast, Fig. 13(b) demonstrates that varying $\beta$ provides an effective route to tune the wave’s frequency while leaving the amplification rate, governed by $\alpha$, nearly unchanged.
In summary, while the amplification rate of waves is primarily determined by $\alpha$, the sign and magnitude of $\beta$ can be used to tune net amplification (through changing the group velocity) and the peak frequency of the wave packet.
Figure 14: Boundary-induced interference of forward and backward wave packets at $|q^{*}|=0.6\pi$ in a 1D lattice with combined nonreciprocal stiffness and gyroscopic damping. (a,e,i) Complex dispersion relations identifying the primary forward ($q^{*}$) and backward ($q^{*}$) components excited by the initial Gaussian packet, alongside the secondary wavenumber $q_{2}$ generated upon boundary reflection. (b,f,j) Ray diagrams illustrating the kinematic trajectories of the direct and reflected propagation paths. (c,g,k) Time-normalized velocity evolution, illustrating the asymmetric amplification and the resulting interference between incident and reflected packets. (d,h,l) Velocity spectra measured at site $n=80$, displaying single or dual frequency peaks depending on the damping parameter $\beta$. Key spectral peaks: $\beta=0$: $(q^{*},\omega^{*})=(0.6\pi,24.38)$; $\beta=-0.2$: primary $q_{1}=0.6\pi$ ($22.89$) and reflected $q_{2}=0.735\pi$ ($25.63$); $\beta=+0.2$: primary $q_{1}=0.6\pi$ ($25.86$) and reflected $q_{2}=0.481\pi$ ($23.00$).
Wave interference through boundary reflection
@@
In a finite system, boundaries are not merely passive reflectors but can be actively utilized to manipulate wave properties. Here, we investigate how boundary reflection can induce wave interference, providing an additional mechanism for tuning the amplified signal’s frequency content. In this context, interference arises from the superposition of the primary forward-propagating wave and the reflected backward-propagating wave, which acquire distinct oscillation frequencies due to the asymmetric dispersion. We achieve this by launching a wave packet near one end of the lattice. The backward-propagating (attenuated) component travels to the opposite boundary, reflects, and then propagates forward, interfering with the original forward-propagating (amplified) component. As we will show, the nonreciprocal damping parameter $\beta$ is crucial in controlling this interference. Figure 14 compares three representative cases: $\beta=0$ (top row), $\beta=-0.2$ (middle row), and $\beta=+0.2$ (bottom row), while keeping the stiffness asymmetry fixed at $\alpha=0.1$.
For the case without gyroscopic damping ($\beta=0$), the dynamics are governed solely by the stiffness asymmetry $\alpha$. The dispersion and ray diagrams [Figs. 14(a–b)] show that the forward and backward components originate from wavenumbers $\pm q^{*}$. Upon reflection, the backward pulse conserves its frequency but scatters into a forward mode with wavenumber $q^{*}$. Consequently, the reflected packet inherits the exact group velocity and growth rate of the primary forward packet. This identical amplification is evident in the time-normalized waterfall plot [Fig. 14(c)], where the relative amplitudes of the two pulses (forward and reflected) remain constant. Accordingly, the velocity spectrum [Fig. 14(d)] at the probe site $n=80$ exhibits a single peak centered around $\omega^{*}$.
For $\beta=-0.2$, the initial wave packet splits into components at $q_{1}$ and $q_{1}$ with distinct frequencies. As shown in Fig. 14(e), the backward pulse oscillates at a higher frequency than the forward one ($\mathrm{Re}[\omega(-q_{1})]>\mathrm{Re}[\omega(+q_{1})]$). After reflecting at the boundary, this high-frequency component couples to a forward mode with positive wavenumber $q_{2}$. Crucially, this mode ($q_{2}$) corresponds to a slower group velocity and a lower growth rate compared to the primary forward packet ($q_{1}$). The ray diagram [Fig. 14(f)] illustrates the resulting divergent trajectories, which prevent the reflected packet from overlapping with the primary pulse. This is confirmed by the waterfall plot [Fig. 14(g)], where the reflected wave lags behind and its relative amplitude decays. Consequently, the spectrum at $n=80$ [Fig. 14(h)] displays two peaks: a dominant one at $\omega_{1}$ (primary wave) and a minor one at $\omega_{2}$ (reflected wave).
Finally, for $\beta=+0.2$, the initial wave packet generates forward and backward components at $q_{1}$ and $q_{1}$, but the dispersion diagram now indicates that the backward pulse oscillates at a lower frequency ($\mathrm{Re}[\omega(-q_{1})]<\mathrm{Re}[\omega(+q_{1})]$) [Fig. 14(i)]. Upon boundary reflection, this lower-frequency pulse scatters into a forward mode at wavenumber $q_{2}$, which possesses a higher group velocity and a larger temporal growth rate than the primary forward wave at $q_{1}$. The ray diagram [Fig. 14(j)] reveals converging trajectories, where the faster reflected packet catches up to the slower primary packet, facilitating spatial overlap. Because the reflected mode grows faster than the primary forward mode, its relative amplitude strengthens more rapidly as the two packets approach each other. This combined effect of convergence and differential growth is directly visible in the waterfall plot [Fig. 14(k)], where the reflected component becomes increasingly dominant in the overlap region. The resulting FFT at $n=80$ [Fig. 14(l)] exhibits two prominent peaks at $\omega_{1}$ and $\omega_{2}$. Notably, the reflected-mode peak ($\omega_{2}$) is substantially stronger here than in the $\beta=-0.2$ case due to its enhanced growth rate. Thus, the combination of boundary reflection and damping asymmetry produces a robust, tunable two-frequency interference response.
@@
The value of $\beta$ along with the boundary thus acts as a switch, controlling which frequencies are mixed and their relative dominance. Therefore, these results demonstrate that boundaries in nonreciprocal lattices can be seen as functional elements that enable complex wave-shaping phenomena in addition to wave amplification.